OpenStudy (anonymous):

Ok.\[f(x)=\left(\begin{matrix}ke ^{-2x} \\ 0\end{matrix}\right)\] \[x \ge 0\]otherwise, Find k

OpenStudy (ash2326):

Is this a probability density function?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Is it piece wise functions?

OpenStudy (anonymous):

Just probability density.

OpenStudy (ash2326):

integrate it between 0 to infinity, it's value is 1, you'll be able to find k

OpenStudy (anonymous):

I'm not sure how to integrate it from infinity... that's the problem I'm having.

OpenStudy (ash2326):

we know for a probability density function \[\int_{-\infty}^{\infty} f(x)=1\] now f(x)=ke^-2x for x\(\ge\)0 \[\int_{0}^{\infty} f(x)=1\] so our integral becomes now f(x)=ke^-2x now \[\int_{0}^{\infty} ke^{-2x}=1\] integral of e^x=e^x so \[(\frac{-1}{2}*k*e^{-2x})=1\] now insert the limits e^(-2x) as x--->\(\infty\) is 0 so we get \[(0-(-1/2*k*e^0)=1\] or \[k/2=1\] or \[\large{k=2}\]

OpenStudy (ash2326):

did you get it order?

OpenStudy (anonymous):

Yes... Just one question, How did (0-(-1/2*k*e^0)=1 become 2?

OpenStudy (ash2326):

see we have, (0-(-1/2*k*e^0)=1 this is an equation now e^0=1 so (0+1/2k*1)=1 so k/2=1 or k=2

OpenStudy (anonymous):

Ah OK. I just figured that out :) Thanks!

OpenStudy (anonymous):

A computer ink cartridge has a life of X hours. The variable X is modelled by the probability density function \[f(x)=\left(\begin{matrix}kx^{-2} \\ 0\end{matrix}\right) \]\[x \ge 0 -otherwise -\] (a) Find K